Global Toys Inc., imposes a payback cutoff of three years for its international investment projects. Assume the company has the following two projects available.



Year Cash Flow (A) Cash Flow (B)
0
Answer #1

media%2F144%2F14469186-c1d3-45b8-9626-85

for project A the payback period is 2.37 years

for project the payback period is 3.08 years
Answer #2

pay back period for A =(1+1+9000/24000) =2.375 years or 2 years and four and a half months


project B = (1+1+1+33000/260000) = 3.127 years or 3 years and 1 and half month (appx)